OpenStudy (anonymous):

What is the slope of the line ? y=-5x-9

OpenStudy (ashes.boom):

in the equation y=mx+b m = slope b = y intercept so in this equation the -5 is your slope

OpenStudy (anonymous):

Slope intercept form of and equation: y=mx+b m=slope b=y int What do you think the slope of your equation is? =)

OpenStudy (anonymous):

Thanks youu ,, What is the y-intercept of the line ? y=-2x-8

OpenStudy (anonymous):

-8 is your y int =)

OpenStudy (ashes.boom):

@CountryGurl15 is right :)

OpenStudy (anonymous):

lol yes thx @Ashes.boom =)

OpenStudy (ashes.boom):

No prob :P lol

OpenStudy (anonymous):

What is the slope, and what is the y-intercept of the line ? y=5/4x+9 Slope ?/? y-Intercept ?

OpenStudy (anonymous):

remember slope is the mx part and the y int is the b part, so what is the slope and the y int? =)

OpenStudy (anonymous):

Y int is 9 i think

OpenStudy (anonymous):

5/4 is slope ?

OpenStudy (ashes.boom):

Yep :) Good job

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

What is the slope-intercept form of this equation? 8x + 6y = 12

OpenStudy (ashes.boom):

okay so slope intercept form is the y=mx+b that we keep bringing up so what you want to do is get y on one side of the equation and the rest on the other side so do you know how to do that or do you want help?

OpenStudy (anonymous):

B? y=-4/3x+6 I think

OpenStudy (ashes.boom):

almost the last term is a little off

OpenStudy (anonymous):

A? y=-4/3x+2

OpenStudy (ashes.boom):

yes thats right :)

OpenStudy (anonymous):

file:///C:/Users/Student/Downloads/School.jpg

OpenStudy (anonymous):

Lol thank you

OpenStudy (anonymous):

Score: 5/5 points Status: Mastered Date Taken: Wed, Apr. 8, 2015 12:14 PM EDT

OpenStudy (ashes.boom):

Nice Job!!